OpenStudy (anonymous):

How do you do this? @phi @jdoe0001

OpenStudy (jdoe0001):

i = prt solve for "r" you're given pretty much everyone else keep in mind that the current balance is 2,300.37 and the original amount was 874 so " i " or the interest earned on it, is the difference between the current balance and the principal

OpenStudy (anonymous):

I'm confused, which is p 2,300... or 874?

OpenStudy (jdoe0001):

\(\bf P=\textit{original amount deposited}\to \$p\\ r=rate\to r\%\to \frac{r}{100}\\ t=years\)

OpenStudy (anonymous):

2,300.37* r* 874= 1/29716r

OpenStudy (anonymous):

@jdoe0001

OpenStudy (jdoe0001):

well.. you want to solve for "r" first in i = prt and then plug in the values

OpenStudy (anonymous):

uhh.. how can you solve for r without plugging in values?

OpenStudy (jdoe0001):

well... you can... but once you have it solve for "r", THEN you plug in the values :)

OpenStudy (jdoe0001):

well.. you can't I meant.. but anyhow,.. you plug in the values after. thus

OpenStudy (anonymous):

Please walk me through this, I tried and got what I showed you above.

OpenStudy (jdoe0001):

well... if you solve for "r"... what do you get?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

5.974472e10

OpenStudy (jdoe0001):

well... say \(\bf i = p{\color{blue}{ r}}t\implies {\color{blue}{ r}}=\square ?\) in variables that is and then you plug in the values for "i", "p", and "t"

OpenStudy (phi):

It helps to understand the "big picture" you give some money to a bank. we call the amount p (short for principal) at some time in the future (t years later), they give back to you more money than you started with.

OpenStudy (phi):

In other words, they return the amount you deposited ( that is, p) plus an additional amount (called the interest), and "i" for short

OpenStudy (anonymous):

:DD This makes a lot more sense now!

OpenStudy (anonymous):

I solved for r = 1.014

OpenStudy (phi):

if you start with 874, what do we call that ?

OpenStudy (phi):

874 is p if they return 2300.37, that is the original 874 plus the interest. any idea what amount the interest is?

OpenStudy (anonymous):

the difference between 874 and 2300.37 or the sum ?

OpenStudy (phi):

if you get confused, use small numbers. say you put in $1 and a year later you get back $3. How much of that is interest?

OpenStudy (anonymous):

$2

OpenStudy (phi):

and how did you get 2? you did 3 - 1 in other words, final amount - original amount use that same idea with 874 and 2300.37

OpenStudy (anonymous):

I did, 1426.37 . I set it up like this -> 874*r*34 = 1426. 37 r= 1.0145

OpenStudy (phi):

874*r*34 = 1426. 37 that is good you can divide both sides by 34 then divide both sides by 874

OpenStudy (phi):

or you could multiply 874*34 to get 29716 29716 r = 1426.37 and now divide both sides by 29716

OpenStudy (anonymous):

^See that's all I needed someone to tell me !!!!! >.<

OpenStudy (anonymous):

0.048

OpenStudy (anonymous):

4.8% or are there more steps?

OpenStudy (phi):

that is the answer!

OpenStudy (anonymous):

Thank you sooo much; you helped me in such an understandable and easy way!